The total area of a regular tetrahedron is 110.85 m2, determine its base edge in m.
1
Expert's answer
2020-12-08T07:39:55-0500

Using the formula

"Area(A)=\sqrt3\times a^2"

where "a" is the length of the side and making "a" the subject of the formula, we obtain:


"a=\frac{1}{3}\times3^{\frac{3}{4}}\times\sqrt A"


Substituting for A,


"a=\frac{1}{3}\times3^{\frac{3}{4}}\times\sqrt{ 110.85}"

therefore:

"a=7.99995m"

Which is approximately "a= 8m"